how do I solve
m = -5/7,(4,2)

Answer by nerdybill(7384)

how do I solve
m = -5/7,(4,2)
.
You don't really "solve" it... rather, you can derive the equation for the line where:
m (slope) is -5/7
and it crosses point (4,2)
.
plug given info into the "point-slope" form:
y - y1 = m(x - x1)
y - 2 = (-5/7)(x - 4)
y - 2 = (-5/7)x + 20/7
y = (-5/7)x + 20/7 + 2
y = (-5/7)x + 20/7 + 14/7
y = (-5/7)x + 34/7 (this is what they're looking for)